OpenStudy (anonymous):

so 2133 base 4 to an equivalent number in base 10 = 183(10)

OpenStudy (anonymous):

nope correct myself 159 is the answer

OpenStudy (anonymous):

\[3+3\times 4+1\times 16+2\times 64\]whatever that is

OpenStudy (anonymous):

huh??

OpenStudy (loser66):

yep I am with you , sorry, 159

OpenStudy (anonymous):

thanks

OpenStudy (loser66):

yw

OpenStudy (anonymous):

is it clear how to get the answer?

OpenStudy (loser66):

@Mandy86 he asks you, not me

OpenStudy (anonymous):

yes i know i was working on another question

OpenStudy (anonymous):

whats clear?

OpenStudy (loser66):

I think he asks you know how to get the answer? the backward process is harder than this forward, I mean convert from base 10 to base.... whatever, is harder than convert from other base to base10 hihi..

OpenStudy (anonymous):

oh sorry i do polynomial form

OpenStudy (loser66):

don't get what you mean, need help? if not, I am out now

OpenStudy (anonymous):

no i got it thank you

OpenStudy (anonymous):

good night